OpenStudy (anonymous):

how to prove 4 points coplaner in vector ex prove for following p(-1,4,3) q(3,2,5) r(-3,8,5) & s (-3,2,1)

OpenStudy (anonymous):

take three vectors as x=q-p y=r-p z=s-p now if the scalar triple product of x, y, z ie., x.(yxz) = 0, the four points are coplanar

OpenStudy (anonymous):

4 points are co-planar if they lie on the same plane note taht a plane is formed only when 2 axis intersect each other if u know 3-d geometry visualise and do in vector notation as i said before u will get some (3,4,1) or 3i+4j+1k if a point is lying on a particular plane then it should have only 2 co-ordinates if it has 3 co-rdiantes then it is said to lie in space(suspended) visualise